# Deciphering Textual Authenticity: A Generalized Strategy through the Lens of Large Language Semantics for Detecting Human vs. Machine-Generated Text

### v1

@@ Abstract
With the recent proliferation of Large Language Models (LLMs), there has been an increasing demand for tools to detect machine-generated text. The effective detection of machine-generated text face two pertinent problems: First, they are severely limited in generalizing against real-world scenarios, where machine-generated text is produced by a variety of generators, including but not limited to GPT-4 and Dolly, and spans diverse domains, ranging from academic manuscripts to social media posts. Second, existing detection methodologies treat texts produced by LLMs through a restrictive binary classification lens, neglecting the nuanced diversity of artifacts generated by different LLMs. In this work, we undertake a systematic study on the detection of machine-generated text in real-world scenarios. We first study the effectiveness of state-of-the-art approaches and find that they are severely limited against text produced by diverse generators and domains in the real world. Furthermore, t-SNE visualizations of the embeddings from a pretrained LLM’s encoder show that they cannot reliably distinguish between human and machine-generated text. Based on our findings, we introduce a novel system, T5LLMCipher, for detecting machine-generated text using a pretrained T5 encoder combined with LLM embedding sub-clustering to address the text produced by diverse generators and domains in the real world. We evaluate our approach across 9 machine-generated text systems and 9 domains and find that our approach provides state-of-the-art generalization ability, with an average increase in F1 score on machine-generated text of 19.6% on unseen generators and domains compared to the top performing existing approaches and correctly attributes the generator of text with an accuracy of 93.6%. We make the code for our proposed approach publicly available at https://anonymous.4open.science/r/LLM-Cipher-84FE/README.md
@@ 1 Introduction
  • New Detection System. We present a new machine-generated text detection system, T5LLMCipher, to robustly attribute texts to source generators using the embeddings from LLM encoders.
  • Comprehensive Evaluation. We demonstrate that classifiers trained for the multi-class classification task of generator attribution show superior generalization and adversarial robustness to binary classifiers in detecting in-the-wild machine-generated text.
@@ 4 Motivation
GLTR-BERT GLTR-GPT2 RoBERTa-D
ChatGPT arxiv -79.4 -100 -74.2
peeread -83.3 -100 -90.0
reddit -28.4 -100 -86.3
wikihow -37.3 -100 -76.7
wikipedia -10.2 -100 -79.4
Dolly arxiv -77.8 -100 -62.4
peeread -87.5 NA -71.9
reddit -52.8 -100 -63.8
wikihow -66.7 -100 -71.4
wikipedia -38.9 -100 -61.5
Table 2: $\Delta$R of existing detectors against black-box adversarial attacks. "+" indicates an improvement in performance, and "-" indicates a deterioration in performance. "NA" values indicate that the baseline recall was 0.
To further show the vulnerability of existing methods, we evaluate the same detectors except for one against a recently introduced adversarial attack, called DFTFooler [41], where we generate adversarial samples designed to evade model detection. We withhold the BERT-D model in this test due to its poor performance shown in Table 1, where there is little room for adversarial attacks to reduce detection performance. Following the evaluation strategy of DFTFooler, we calculate the percentage change in recall, $\Delta$Recall ($\Delta$R). We use this metric to evaluate model resilience to adversarial attacks because it quantifies the change in how well a model correctly identifies the machine-generated text under these attacks. A significant drop in recall indicates the model’s vulnerability, as it fails to identify many adversarially altered instances correctly. The results presented in Table 2 underscore the vulnerabilities of current text detection methods against adversarial attacks. Examining the $\Delta$R for both ChatGPT and Dolly generators, we observe a notable trend: GLTR-GPT2 consistently yielded a 100% drop in recall, highlighting its susceptibility to these attacks irrespective of the text domain. Furthermore, while the performance of detectors such as GLTR-BERT and RoBERTa-D varied based on the dataset, several instances showcased deterioration in performance, ranging from 70-90%. While the ($\Delta$R) on BERT-D shows positive values, meaning the adversarial attack actually increased the detection performance on adversarially perturbed samples, the baseline recall was already so low that the perturbations flipped the decisions on some of the originally incorrect predictions to being correct predictions. Such high drops in recall by the models accentuate the pressing need for enhancing the robustness of current detection methods, as even these advanced models demonstrate significant vulnerabilities in the face of adversarial attacks.
@@
In Figure 1, we present the t-SNE visualization of text embeddings derived from wikipedia text comprising both machine and human-generated text, as detailed in the study by Wang et al.  [54]. The preliminary results from the t-SNE experiment are encouraging, showing some separation between human and machine-generated text in the reduced-dimension space created by t-SNE. While there is considerable overlap between the human and machine-generated texts, such initial findings warrant further exploration and analysis into these embeddings. Given that this embedding was never trained to discern human text from machine text, it shows potential to be developed into a robust classifier. The observed separation provides a foundation on which we can apply techniques like contrastive learning or deep neural networks to further refine the distinction between machine-generated and human-generated texts. In the subsequent Figure 2, we delve deeper, showcasing the t-SNE visualizations encompassing six distinct classes: human text and five different machine text generators. Here, the interplay of the classes paints a more complex picture. Figure 2 shows the increased complexity in distinguishing human-generated text from a diverse array of machine-generated sources, as compared to differentiating human text from a singular machine source. However, an interesting observation arises: the distribution of different machine generators in this t-SNE visualization of the LLM embedding suggests that, with further refinement and training, the LLM embedding could be harnessed to attribute the source generator of a specific text piece. This insight broadens the scope of our research, indicating that not only can we detect machine-generated content, but we might also be able to pinpoint its origin.
5 Data
This work leverages diverse datasets collected by Wang et al.  [54] to construct a robust detector for machine-generated text. We focus on the English corpora they released so that we can compare our approach against other in-the-wild machine-generated text datasets. We analyze five sources of human-generated data and five LLMs used to generate machine-generated text from Wang et al.  [54]. These LLMs are ChatGPT (gpt3.5-turbo)  [6], davinci-003  [6], Cohere  [3], Dolly-v2  [4], and BLOOMz  [35]. The human text sources and the corresponding text generation strategies are described below:
  • Reddit ELI5: A collection of 3,000 preprocessed question and response pairs from subreddits r/explainlikeimfive, r/askscience, and r/AskHistorians. Answers to questions are used as prompts for the LLM.
@@ 6 Approach
To effectively discern machine-generated from human-generated text, we leverage the power and sophistication of Large Language Model (LLM) encoders. These encoders, which have been proven adept at generating dense and semantically rich embeddings of textual data [55], serve as the foundation of our methodology. By feeding both human and machine-generated texts into such an encoder, we extract embeddings that inherently capture the unique characteristics and intricacies of each text type. These embeddings are then stored for further analysis. Atop these embeddings, we construct classifiers — from Multilayer Perceptrons (MLPs) and K-Nearest Neighbors (KNN) to more advanced, contrastive learning-based KNN classifiers. We call this approach T5LLMCipher, based on the T5 [44] LLM encoder that we use in our implementation. In the ensuing sections, we describe LLM encoders, elucidate why they stand out as promising tools for machine-generated text detection, and provide a detailed exposition on the classifiers designed atop these embeddings. The overall proposed system architecture is shown in Figure 3.
@@ 6.1 LLM Text Encoder
Building on the rich embeddings provided by the LLM encoders, our methodology uses a range of classifiers, each designed to utilize the distinct characteristics of these embeddings. The objective is clear: to leverage the nuances captured within the embeddings to differentiate between human-generated and machine-generated texts.
@@ 6.2.2 K-Nearest Neighbors Classification
We also use the K-Nearest Neighbors (KNN) classifier to distinguish between human and machine-generated texts. At its core, the KNN algorithm attributes a classification label to an unseen sample based on the predominant label of its $k$ closest neighbors. One of the innate strengths of the KNN algorithm is its resistance to overfitting, an issue that often plagues more complex classifiers. Models that are highly susceptible to overfitting excel on their training data but display a marked decline in performance on novel datasets [12]. Several advanced classifiers, despite their sophistication, are known to fall into this trap [16]. The non-parametric nature of KNN circumvents this challenge by drawing on broader patterns and trends present in the data. We carefully choose the training data for our models to be paired in nature, where every human-generated text is coupled with a corresponding machine-generated text on the same topic. This structure is particularly beneficial for KNN. When faced with paired texts that share thematic content but differ in their origins, KNN can sharply focus on the subtle distinctions between them. During inference, the classification decision should predominantly reflect the origin of the text rather than its content.
@@ 7.3 Text Generator Identification
Table 3 is a confusion matrix that details the number of times a given text source was classified as each possible source. The diagonal of the matrix represents true positives or instances where the classifier correctly identified the source. For example, the classifier correctly identified Bloomz-generated text 1238 times and misclassified it 23 times. Notably, Human-generated text seems distinctly characterized, as the classifier correctly classified it 6620 times with a minimal number of misclassifications.
Table 4 delves into the individual performance metrics of the classifier for each text generator. It lists precision, recall, F1-Score, and support for each class. This table shows that the classifier performs exceptionally well in identifying Human and Bloomz text, with F1 scores of 0.988 and 0.977, respectively. ChatGPT and Cohere texts are fairly distinguishable with F1-Scores above 0.870, whereas Davinci and Dolly have the lowest F1-Scores, suggesting these might be harder to distinguish or have characteristics that overlap with other classes. Additionally, we visualize the model’s proficiency in differentiating between machine-generated and human-generated texts through the t-SNE presented in Figure 4. For this visualization, we extracted features from the layer before the classification layer, which has a vector size of 256. The domains showcased include Arxiv and Wikihow, spanning five distinct generators. In this figure, human and machine samples manifest as more defined clusters than the distribution observed in Figure 1. We further visualize the model’s ability to attribute text to source generators in each domain in Figure 5. There we show the t-SNE visualization of 6 class classifier embeddings on each of the different generator’s texts and see that each one forms a distinct cluster with minimal overlap. Compared to the original embedding space shown in Figure 2, we now see excellent separation between classes. Overall, these experiments show that machine-generated that the source generator of texts is feasible, and that different generators have unique identifiable characteristics, or "fingerprints," within their generated texts.
@@ 7.6 In-the-wild Experiment
AI Writer Article Forge Kafkai Reddit Bot
T5LLMCipher-MC 62.3 84.4 61.8 84.6
T5LLMCipher-Bi 59.2 78.8 46.2 82.1
T5LLMCipher-C-KNN 57.2 83.2 49.2 82.9
T5LLMCipher-KNN 37.2 45.4 32.8 67.0
RoBERTaLLMCipher-MC 61.5 81.2 41.9 81.8
@@
We train our models on the five generators across five domains from the M4 dataset. The 6 class T5LLMCipher-MC was trained to attribute text to Human class and each of the five machine-text generators. For the testing of the 6 class T5LLMCipher-MC classifier in the cross-domain experiment, if the classifier detects a sample as a human sample, it is treated as a human sample, and if it is detected as any of the five machine generators, it is treated as a machine sample. All other classifiers were trained and tested on the binary classification task of distinguishing between machine and human text. Furthermore, to demonstrate the impact of the backbone LLM encoder, we also test a 6 class MLP classifier using RoBERTa instead of T5, which is indicated as RoBERTaLLMCipher-MC.
We find that T5LLMCipher-MC provides the best results with an average F1 score of 73.3% across the four in-the-wild dataset. We also see that our models have the toughest time generalizing to the Kafkai generations, and the easiest time detecting RedditBot generations. We also find that T5LLMCipher-MC with T5 as the LLM encoder outperforms RoBERTaLLMCipher-MC with RoBERTa as the LLM encoder.

### v3

@@ Abstract
With the recent proliferation of Large Language Models (LLMs), there has been an increasing demand for tools to detect machine-generated text. The effective detection of machine-generated text face two pertinent problems: First, they are severely limited in generalizing against real-world scenarios, where machine-generated text is produced by a variety of generators and spans diverse domains. Second, existing detection methodologies treat texts produced by LLMs through a restrictive binary classification lens, neglecting the nuanced diversity of artifacts generated by different LLMs, each of which exhibits distinctive stylistic and structural elements. In this work, we undertake a systematic study on the detection of machine-generated text in real-world scenarios. We first study the effectiveness of state-of-the-art approaches and find that they are severely limited against text produced by diverse generators and domains in the real world. Furthermore, t-SNE visualizations of the embeddings from a pretrained LLM’s encoder show that they cannot reliably distinguish between human and machine-generated text. Based on our findings, we introduce a novel system, T5LLMCipher, for detecting machine-generated text using a pretrained T5 encoder combined with LLM embedding sub-clustering to address the text produced by diverse generators and domains in the real world. We evaluate our approach across 9 machine-generated text systems and 9 domains and find that our approach provides state-of-the-art generalization ability, with an average increase in F1 score on machine-generated text of 11.9% on unseen generators and domains compared to the top performing supervised learning approaches and correctly attributes the generator of text with an accuracy of 93.6%. We make the code for our proposed approach publicly available at https://github.com/SecureAIAutonomyLab/LLM-Cipher
@@ 1 Introduction
  • New Detection System. We present a new machine-generated text detection system, T5LLMCipher, to robustly detect text as being machine generated using the embeddings from LLM encoders.
  • Comprehensive Evaluation. We demonstrate that classifiers trained for the multi-class classification task of generator attribution show superior generalization and adversarial robustness to binary supervised classifiers in detecting in-the-wild machine-generated text.
@@ 3.2 Threat Model
4 Data
This work leverages diverse datasets collected by Wang et al.  [61] in the M4 dataset to construct a robust detector for machine-generated text. We focus on the English corpora they released so that we can compare our approach against other in-the-wild machine-generated text datasets. We analyze five sources of human-generated data and five LLMs used to generate machine-generated text from the M4 dataset [61], for a total of 126,950 text samples. These LLMs are ChatGPT (gpt3.5-turbo)  [6], davinci-003  [6], Cohere  [3], Dolly-v2  [4], and BLOOMz  [41]. Additional discussions on dataset assumptions and experiments that assess the quality of the machine-generated texts in the are detailed in Appendix A.2. This appendix presents results from an automatic evaluation using Universal Sentence Encoder [20], alongside human evaluations that assess the stylistic and semantic similarity between the human and machine-generated texts. The human text sources and the corresponding text generation strategies are described below, with further prompt details provided in Appendix A.1:
  • Reddit ELI5: A collection of 3,000 preprocessed question and response pairs from subreddits r/explainlikeimfive, r/askscience, and r/AskHistorians. Questions are used as prompts for the LLM to generate an answer.
@@ 5 Motivation
BERT-D GLTR-BERT GLTR-GPT2 RoBERTa-D
ChatGPT arxiv +21.4 -79.4 -100 -74.2
peeread NA -83.3 -100 -90.0
reddit NA -28.4 -100 -86.3
wikihow NA -37.3 -100 -76.7
wikipedia NA -10.2 -100 -79.4
Dolly arxiv +27.3 -77.8 -100 -62.4
peeread -4.9 -87.5 NA -71.9
reddit +23.5 -52.8 -100 -63.8
wikihow +10.3 -66.7 -100 -71.4
wikipedia -0.0 -38.9 -100 -61.5
Table 2: $\Delta$R of existing detectors against black-box adversarial attacks. "+" indicates an improvement in performance, "-" indicates a deterioration in performance. "NA" values indicate that the baseline recall was 0.
To further show the vulnerability of existing methods, we evaluate the same detectors except for one against a recently introduced adversarial attack, called DFTFooler [47], where we generate adversarial samples designed to evade model detection. Following the evaluation strategy of DFTFooler, we calculate the percentage change in recall, $\Delta$Recall ($\Delta$R). We use this metric to evaluate model resilience to adversarial attacks because it quantifies the change in how well a model correctly identifies the machine-generated text under these attacks. A significant drop in recall indicates the model’s vulnerability, as it fails to identify many adversarially altered instances correctly. The results presented in Table 2 underscore the vulnerabilities of current text detection methods against adversarial attacks. Examining the $\Delta$R for both ChatGPT and Dolly generators, we observe a notable trend: GLTR-GPT2 consistently yielded a 100% drop in recall, highlighting its susceptibility to these attacks irrespective of the text domain. Furthermore, while the performance of detectors such as GLTR-BERT and RoBERTa-D varied based on the dataset, several instances showcased deterioration in performance, ranging from 70-90%. While the ($\Delta$R) on BERT-D shows positive values, meaning the adversarial attack actually increased the detection performance on adversarially perturbed samples, the baseline recall was already so low that the perturbations flipped the decisions on some of the originally incorrect predictions to being correct predictions. Such high drops in recall by the models accentuate the pressing need for enhancing the robustness of current detection methods, as even these advanced models demonstrate significant vulnerabilities in the face of adversarial attacks.
@@
In Figure 1, we present the t-SNE visualization of text embeddings derived from wikipedia text comprising both machine and human-generated text, as detailed in the study by Wang et al.  [61]. The preliminary results from the t-SNE experiment are encouraging, showing some separation between human and machine-generated text in the reduced-dimension space created by t-SNE. While there is considerable overlap between the human and machine-generated texts, such initial findings warrant further exploration and analysis into these embeddings. Given that this embedding was never trained to discern human text from machine text, it shows potential to be developed into a robust classifier. The observed separation provides a foundation on which we can apply techniques like contrastive learning or deep neural networks to further refine the distinction between machine-generated and human-generated texts. In the subsequent Figure 2, we delve deeper, showcasing the t-SNE visualizations encompassing six distinct classes: human text and five different machine text generators. Here, the interplay of the classes paints a more complex picture. Figure 2 shows the increased complexity in distinguishing human-generated text from a diverse array of machine-generated sources, as compared to differentiating human text from a singular machine source. Given that there is some marginal difference in the distribution of the embeddings between the different generators even in this unsupervised visualization suggests that, with further refinement and training, the LLM embedding could be harnessed to attribute the source generator of a specific text piece. This insight broadens the scope of our research, indicating that not only can we detect machine-generated content, but we might also be able to pinpoint its origin.
@@ 6 Approach
To effectively discern machine-generated from human-generated text, we leverage the power and sophistication of Large Language Model (LLM) encoders. These encoders, which have been proven adept at generating dense and semantically rich embeddings of textual data [62], serve as the foundation of our methodology. By feeding both human and machine-generated texts into such an encoder, we extract embeddings that inherently capture the unique characteristics and intricacies of each text type. These embeddings are then stored for further analysis. Atop these embeddings, we construct classifiers from Multilayer Perceptrons (MLPs) and K-Nearest Neighbors (KNN) to more advanced, contrastive learning-based KNN classifiers. We call this approach T5LLMCipher, based on the T5 [50] LLM encoder that we use in our implementation. In the ensuing sections, we describe LLM encoders, elucidate why they stand out as promising tools for machine-generated text detection, and provide a detailed exposition on the classifiers designed atop these embeddings. The overall proposed system architecture is shown in Figure 3.
@@ 6.1 LLM Text Encoder
Building on the rich embeddings provided by the LLM encoders, our methodology uses a range of classifiers, each designed to utilize the distinct characteristics of these embeddings. The following subsubsections outline the different types of Cipher models that we use to build off of these embeddings.
@@ 6.2.2 K-Nearest Neighbors Classification
We also use the K-Nearest Neighbors (KNN) classifier to distinguish between human and machine-generated texts. At its core, the KNN algorithm attributes a classification label to an unseen sample based on the predominant label of its $k$ closest neighbors. One of the innate strengths of the KNN algorithm is its resistance to overfitting, an issue that often plagues more complex classifiers. Models that are highly susceptible to overfitting excel on their training data but display a marked decline in performance on novel datasets [12]. Several advanced classifiers, despite their sophistication, are known to fall into this trap [17]. The non-parametric nature of KNN circumvents this challenge by drawing on broader patterns and trends present in the data. We carefully choose the training data for our models to be paired in nature, where every human-generated text is coupled with a corresponding machine-generated text on the same topic. In the training datasets that are used for our models, each sample is a pair consisting of human-written text and the machine-generated text that was generated using this human-written text. This structure is particularly beneficial for KNN. When faced with paired texts that share thematic content but differ in their origins, KNN can sharply focus on the subtle distinctions between them. During inference, the classification decision should predominantly reflect the origin of the text rather than its content.
@@ 7.3 Text Generator Identification
Table 3 is a confusion matrix that details the number of times a given text source was classified as each possible source. The diagonal of the matrix represents true positives or instances where the classifier correctly identified the source. For example, the classifier correctly identified Bloomz-generated text 98.1% of the time. Human-generated text also seems to be distinctly characterized, as the classifier correctly attributed human texts to being human-generated 98.2% of the time.
Table 4 delves into the individual performance metrics of the classifier for each text generator. It lists precision, recall, F1-Score, and support for each class. This table shows that the classifier performs exceptionally well in identifying Human and Bloomz text, with F1 scores of 0.988 and 0.977, respectively. ChatGPT and Cohere texts are fairly distinguishable with F1-Scores above 0.870, whereas Davinci and Dolly have the lowest F1-Scores, suggesting these might be harder to distinguish or have characteristics that overlap with other classes. Additionally, we visualize the model’s proficiency in differentiating between machine-generated and human-generated texts through the t-SNE presented in Figure 4. For this visualization, we extracted features from the layer before the classification layer, which has a vector size of 256. The domains showcased include Arxiv and Wikihow, spanning five distinct generators. In this figure, human and machine samples manifest as more defined clusters than the distribution observed in Figure 1. We further visualize the model’s ability to attribute text to source generators in each domain in Figure 5. There we show the t-SNE visualization of 6 class classifier embeddings on each of the different generator’s texts and see that each one forms a distinct cluster with minimal overlap. Compared to the original embedding space shown in Figure 2, we now see excellent separation between classes. Overall, these experiments show that the identification of the source generator of texts is feasible, and that different generators have unique identifiable characteristics, or "fingerprints," within their generated texts.
@@ 7.6 In-the-wild Experiment
AI Writer Article Forge Kafkai Reddit Bot Avg
T5LLMCipher-MC 62.3 84.4 61.8 84.6 73.3
T5LLMCipher-Bi 59.2 78.8 46.2 82.1 66.6
T5LLMCipher-C-KNN 57.2 83.2 49.2 82.9 68.1
T5LLMCipher-KNN 37.2 45.4 32.8 67.0 45.6
RoBERTaLLMCipher-MC 61.5 81.2 41.9 81.8 66.6
BERT-D 68.6 70.6 55.6 67.3 65.5
GLTR-BERT 49.2 83.8 29.0 73.8 59.0
GLTR-GPT2 29.6 83.1 29.2 83.5 56.4
RoBERTa-D 20.9 48.7 37.6 76.8 46.0
RADAR 50.2 61.6 60.2 18.0 48.0
Fast-DetectGPT 32.1 52.9 98.7 97.5 70.3
@@
We train our models on the five generators across five domains from the M4 dataset. The 6 class T5LLMCipher-MC was trained to attribute text to Human class and each of the five machine-text generators. This model, which was trained to attribute texts to known generators, will instead be used for the binary task of attributing texts as machine-generated or human-generated. For the testing of the 6 class T5LLMCipher-MC classifier in the cross-domain experiment, if the classifier detects a sample as a human sample, it is treated as a human sample, and if it is detected as any of the five machine generators, it is treated as a machine sample. The rationale behind this testing strategy for our T5LLMCipher-MC is that these new samples from unseen LLMs will likely be far away from the human text cluster and may share some characteristics with one of many known generators that are already included in its training. All other classifiers were trained and tested on the binary classification task of distinguishing between machine and human text. Furthermore, to demonstrate the impact of the backbone LLM encoder, we also test a 6 class MLP classifier using RoBERTa instead of T5, which is indicated as RoBERTaLLMCipher-MC. We also compare against the four models discussed in Section 5 (BERT-D, GLTR-BERT, GLTR-GPT2, and RoBERTa-D) [47] by training these models on the same training dataset used to train our classifiers to provide an apples-to-apples comparison against existing supervised detection models. To compare against black-box detection methods, we examine the RADAR model [31], a framework designed for detecting machine-generated texts through adversarial learning. Central to RADAR are two key components: a paraphraser and a detector. The paraphraser, built on the T5-large model [50], is tasked with altering machine-generated text in an effort to bypass detection. Concurrently, the detector, leveraging the RoBERTa-large model [38], is trained to identify both the original AI texts and those that have been paraphrased. This adversarial training loop ensures that RADAR continuously evolves in its ability to discern AI-generated content. The training data for RADAR comprises a human text corpus, sourced from the OpenWebText dataset [28], and machine-generated text, which employs the Vicuna-7B model to create texts based on prefixes extracted from the human-text corpus. We further compare against Fast-DetectGPT [16], another black-box machine-text detection algorithm. This method quantifies the curvature of the conditional probability function, which is indicative of the likelihood of text being machine-generated based on the comparative evaluation of the probabilities of given and alternative token choices. The positive curvature at a point suggests a higher likelihood of the text being machine-generated, whereas a close-to-zero curvature indicates human authorship. GPT-Neo 2.7B [19] was used as the scorer model to obtain token probabilities.
We find that T5LLMCipher-MC provides the best results with an average F1 score of 73.3% across the four in-the-wild datasets. We also observe that our models have the toughest time generalizing to the Kafkai generations, and the easiest time detecting RedditBot generations. We also find that T5LLMCipher-MC with T5 as the LLM encoder outperforms RoBERTaLLMCipher-MC with RoBERTa as the LLM encoder. Some other notable results come from Fast-DetectGPT, where Kafkai and Reddit Bot generations are detected with very high F1 scores. This can be explained by the underlying scoring model used by Fast-DetectGPT, which is a GPT variant. Kafkai and Reddit Bot machine-texts use models from the GPT family. This corroborates with the results from the work of Bao et al. [16] where they show that their detection methodology generalizes well to other machine text generators from the same family. However, their method is shown to struggle in discerning human from machine-generated texts from different families of LLMs, as is shown from the poor detection performance on AI Writer and Article Forge texts.